# How and Why LLMs Use Deprecated APIs in Code Completion? An Empirical Study

### v2

@@ IV-B2 Root Cause Analysis
The contextual characteristics of the input completion prompts can significantly influence LLMs’ use of deprecated APIs. Since the completions were generated based on the input prompts, specific contexts can lead LLMs to use deprecated APIs. As presented above, the LLMs showed significantly different DUR for prompts from $\mathcal{O}$ and $\mathcal{U}$. Upon comparing the prompts from the two datasets, we found that contextual characteristics of the prompts, such as specific objects, control flows, and data flows, often lead the LLMs’ predictions, i.e., whether to use deprecated APIs or their replacements.
@@
Finding 6: The contextual characteristics of the input prompts, such as the defined objects, control flows, and data flows, is a significant influence on the use of deprecated APIs.
IV-C RQ3: Library Perspective for Status Quo and Cause Analysis
@@
We also conducted a detailed analysis to examine the LLMs’ completions across different libraries.
@@ IV-C1 Status Quo Analysis
The status quo from library perspective focuses on how the AUP and DUR metrics vary with different libraries.
@@
Finding 7: Most LLMs exhibited relatively low API usage plausibility across the 8 libraries, with AUP below 30%. CodeLlama-7b, CodeGen-6b, and CodeGen-2b achieved better AUP for Pandas (about 45%-65%) and TensorFlow (around 40%).
Deprecated Usage Rate. The results are presented in the scatter plots shown in Figure 5, where each data point represents the DUR of a particular LLM for a specific library. The results reveal significant differences in the usage of deprecated APIs across the libraries, with DUR ranging from approximately 0% to 100%. Specifically, LLMs generally showed low DUR (around or below 20%) for Numpy, scikit-learn, and TensorFlow. In contrast, LLMs exhibited consistently high DUR for SciPy (approximately 30%-50%) and PyTorch (approximately 50%-70%), and unstable DUR for Pandas (approximately 30%-80%) and Transformers (approximately 40%-100%).
@@
Finding 8: LLMs showed significant differences in the usage of deprecated APIs across various libraries, with DUR ranging from approximately 0% to 100%. LLMs exhibited consistently high DUR for SciPy and PyTorch, and unstable DUR for Pandas and Transformers.
IV-C2 Root Cause Analysis
@@
The characteristics of API deprecations during library evolution also significantly impacted the use of deprecated APIs in LLM-based code completions. Some APIs were deprecated due to simple package refactoring, leading to similar usage patterns for the deprecated APIs and their replacements. For example, in the version 1.9.0 release of PyTorch, many APIs for tensor linear algebra were moved from the torch package to the torch.linalg package without changes to the API parameters or usage patterns (e.g., torch.lstsq $\to$ torch.linalg.lstsq). In such cases, LLMs found it more difficult to distinguish between deprecated APIs and their replacements, resulting in more frequent use of deprecated APIs in the predicted completions.
Finding 9: The characteristics of API deprecations during library evolution significantly impacted the use of deprecated APIs in LLM-based code completions. Minor changes between deprecated APIs and their replacements, such as simple package refactoring, often led to more pronounced issues with deprecated API usages.
V Mitigation Approaches
@@ V-A Motivation
For model training, a direct mitigation method is to clean up the code containing deprecated APIs from the training corpora. However, this is impractical due to the constant evolution of libraries. New API deprecations would necessitate repeatedly rebuilding the training corpora and retraining the model. From the library perspective, we cannot control the evolution of the library and API deprecation. Therefore, mitigation efforts should focus on decoding strategies in model inference and prompt engineering.
@@ V-B Two Lightweight Approaches
where $\mathbb{T}^{I}$ and $\mathbb{T}^{O}$ are the input token sequence and output token sequence, respectively. In the context of LLM-based code completion, $\mathbb{T}^{I}$ corresponds to the input prompt $pmpt$, and $\mathbb{T}^{O}$ corresponds to the predicted completion $comp$ (line 1). When $comp$ contains a deprecated API $dep$ (line 3), we need to perform fixing approaches, i.e., the Fix procedure, to replace $dep$ with the corresponding replacement $rep$ (line 4). Note that during the Contains procedure in line 3, API alias resolution is conducted similarly to the process described in Section III-B1.
@@ 4             $comp\leftarrow\textbf{{Fix}}(pmpt,comp,dep,rep)$
The Fix procedure can be achieved by reconstructing input, through either (i) replacing the deprecated API tokens or (ii) inserting an additional replacing prompt, and then regenerating output:
  • Approach 1 - ReplaceAPI: Replacing Deprecated API Tokens then Regenerating. As shown in Algorithm 2, the ReplaceDep procedure involves removing the tokens corresponding to $dep$ and any subsequent tokens from $comp$, then appending the tokens of $rep$ (line 2). This results in a prefix $prefix$ that includes $rep$. The $prefix$ is concatenated with the $pmpt$, and the LLM generates the suffix (line 3), i.e., arguments for $rep$ and the remaining tokens to complete the code line. This concatenation forms the fixed completion $comp^{*}$.
  • Approach 2 - InsertPrompt: Inserting Additional Replacing Prompt then Regenerating. As shown in Algorithm 3, the CreateDepPmpt procedure constructs an additional replacing prompt $pmpt^{\prime}$ (line 2), formatted as inline comments to guide the LLM to use $rep$ instead of $dep$. By expressing the replacing instruction into inline comments, the re-written prompt (i.e., $pmpt\oplus pmpt^{\prime}$) can be naturally processed by the LLM to continue writing the code. The replacing prompt is structured as follows, where “{comp}”, “{dep}”, and “{rep}” are placeholders for the original completion, deprecated API, and replacing API, respectively, and “...” represents the indentation to ensure syntax correctness.
@@ 4       $comp^{*}\leftarrow prefix\oplus suffix$
Algorithm 2 Approach 1: Replacing Deprecated API Tokens then Regenerating
@@ 3       $comp^{*}\leftarrow\textbf{{LLM}}(pmpt\oplus pmpt^{\prime})$
Algorithm 3 Approach 2: Inserting Additional Replacing Prompt then Regenerating
@@ V-C1 Evaluation Procedure
For each LLM, we selected up-to-dated samples from $\mathcal{U}$ where the LLM predicted bad completions using deprecated APIs, i.e.,
$\mathcal{T}=\{(pmpt,dep\to rep)\in\mathcal{U}:\text{annotate}(\textbf{{LLM}}(pmpt))=\emph{bad}\},$
to constitute the evaluation data. These samples were chosen because they have corresponding ground-truth completions (e.g., the line following the prompt in Figure 2), which are essential for assessing the effectiveness of the proposed fixing approaches.
For each sample, we employed the deprecation-aware code completion illustrated in Algorithm 1 with the two fixing approaches ReplaceAPI and InsertPrompt to prompt the LLM to generate a completion.
We used the following three metrics to assess the effectiveness of the proposed approaches:
  • Fixed Rate (FR): This metric indicates the proportion of good completions predicted by the fixing approaches.
@@ V-C2 Results and Analysis
The evaluation results of the proposed fixing approaches are presented in Table IV.
ReplaceAPI. Using the ReplaceAPI fixing approach (Algorithm 2), all the LLMs achieve high fixed rates (FR), with values exceeding 85%. Failures in fixing are mainly due to syntax errors or incorrect function calls caused by erroneous tokens following the replaced APIs. For example, consider a bad completion: “meta_graph_def=tf.saved_model.loader.load(…)” predicted by the original completion procedure (line 1 of Algorithm 1). ReplaceAPI replaces the deprecated API tf.saved_model.loader.load with its replacement tf.saved_model.load, producing a $prefix$ of “meta_graph_def=tf.saved_model.load” (line 2 of Algorithm 2). However, CodeGen-2b then predicts a $suffix$ of “_meta_graph_def(…)” (line 3 of Algorithm 2), resulting in an erroneous function call: tf.saved_model.load_meta_graph_def(). This issue arises because the replacement operation in ReplaceAPI can disrupt the naturalness of the code context [65] and the LLMs’ decoding process. An additional interesting finding is that for the three versions of CodeGen, the FR decreases as the model size increases. This suggests that larger models might be more sensitive to interventions in the decoding process.
The completions generated by LLMs using the ReplaceAPI approach exhibit high edit similarity (ES), exceeding 80%, and exact match rates between 30% and 50%, with these rates increasing alongside model size. The inaccuracies in the completions often involve incorrect return values and arguments for the replacing APIs. Incorrect return values arise because the replacing API might include different elements compared to the deprecated API, and the ReplaceAPI approach cannot resolve such inconsistencies in the $prefix$ (line 2 of Algorithm 2). The incorrect arguments are primarily due to the LLMs’ limitations in correctly utilizing replacing APIs, especially those with complex argument lists.
InsertPrompt. When applying the InsertPrompt fixing approach (Algorithm 3), the LLMs exhibited significantly varied fixed rates, ranging from 25.7% to 97.2%. This variation suggests that larger models generally possess a stronger capacity to interpret and utilize inserted prompts formatted as inline comments. An exception to this trend is DeepSeek-1.3b, which achieved a notably high FR of 93.5%. This success can be attributed to using the deepseek-coder-1.3b-instruct version, which has robust zero-shot instruction-following capabilities. Moreover, the FR differences among various LLMs highlight their sensitivity to prompt construction [66, 67]. This sensitivity indicates that different LLMs may require specialized additional prompts in InsertPrompt for optimal performance.
TABLE IV: Evaluation Results of Proposed Approaches
@@
Finding 10: The proposed ReplaceAPI effectively addresses deprecated API usage for all open-source LLMs, achieving fix rates exceeding 85%. The fixed completions also demonstrate acceptable accuracy compared to ground-truth completions. While InsertPrompt does not currently achieve sufficient effectiveness and accuracy in fixing completions containing deprecated API usage, it shows potential for future exploration.
@@ VI-A Implications
Validating LLM-Generated Code Completions and Issuing Alerts for Deprecated API Usages. Our evaluation study reveals that LLMs frequently use deprecated APIs during code completion. Such deprecated API usages can be easily overlooked [19], potentially introducing bugs or security vulnerabilities into software projects. Therefore, implementing a validation mechanism for deprecated API usage in LLM-generated code completions is crucial to ensure the reliability of the code. Such a validation mechanism can be further integrated into the pose-processing of LLM-based code completion, such as issuing alerts to developers.
@@
Designing Agent & Multi-Agent Systems for Incorporating Library Evolution into LLM-Driven Software Development. In modern software development driven by LLMs, the issues brought by library evolution are encountered not only in code completion. With advancements in AI agents [77, 78, 79], we can potentially develop autonomous agents or multi-agent systems capable of automatically discovering deprecated API usages, identifying correct replacements, upgrading dependent libraries, and fixing the code. To ensure comprehensive recognition of deprecated API usage and up-to-date fixes, we should design an effective multi-agent collaboration pipeline. One agent should scan the generated code to identify all pieces related to API usage. Another agent should continuously fetch information online, checking the latest official API documentation to aid in discovery and correction. Finally, a dedicated agent should be responsible for implementing the necessary library upgrades and code fixes. Such an agent system can fundamentally address the issues discussed in this article.
@@ VI-B Threats to Validity
Internal Threats. The primary threat to the internal validity of our study is the soundness of the static analysis used for function location and result annotation. Given that Python is a dynamic programming language, the lightweight object type resolution method we employed may have missed some function calls of deprecated and replacing APIs during the matching process. In the future, we aim to address this issue by implementing advanced type inference techniques. Additionally, our study currently focuses on function-level API deprecation, overlooking parameter-level deprecations. Future research should investigate a broader range of deprecated API categories to provide a more comprehensive analysis.
External Threats. A primary threat to the external validity of our study lies in the choice of Python libraries and the evaluated LLMs. To mitigate this threat, we reused libraries examined in previous studies and introduced three popular deep learning libraries to ensure diversity and timeliness. For the LLMs, we selected models covering various architectures, model sizes, training corpora, and training strategies to ensure the generalizability of our findings. Another external threat is that the study was conducted solely on the Python language, which may limit the applicability of our findings to other languages such as Java and C#. In the future, we plan to explore the impact of library evolution on LLM-based code completion across a broader range of programming languages.

### v3

@@ IV-B2 Potential Cause Discussion
The contextual characteristics of the input completion prompts can significantly influence LLMs’ use of deprecated APIs. Since the completions were generated based on the input prompts, specific contexts can lead LLMs to use deprecated APIs. As presented above, the LLMs showed significantly different DUR for prompts from $\mathcal{O}$ and $\mathcal{U}$. Upon comparing the prompts from the two datasets, we found that contextual characteristics of the prompts, such as specific variables and function calls, lead to the discrepancy of LLMs’ predictions, i.e., whether to use deprecated APIs or their replacements.
To quantitatively assess differences in contextual characteristics, we conducted lightweight contextual feature extraction for prompts from $\mathcal{O}$ and $\mathcal{U}$ using static analysis. Specifically, we grouped prompts according to their reference API, denoted as $api$, with each group represented as $\mathcal{P}(api)$. For each prompt in $\mathcal{P}(api)$, we identified the function used to construct it (cf. Section III-B) and extracted the variable and function call names in this function. After analyzing all prompts within $\mathcal{P}(api)$, the extracted variable and function call names were compiled into a contextual feature set, represented by $\mathbf{Feat}(\mathcal{P}(api))$, for the prompt group. Through this approach, for each API mapping $dep\to rep$, we obtained two paired prompt groups $\mathcal{P}(dep)\subset\mathcal{O}$ and $\mathcal{P}(rep)\subset\mathcal{U}$, along with their feature sets, $\mathbf{Feat}(\mathcal{P}(dep))$ and $\mathbf{Feat}(\mathcal{P}(rep))$. We then assessed the contextual similarity between $\mathcal{P}(dep)$ and $\mathcal{P}(rep)$ using $\frac{|\mathbf{Feat}(\mathcal{P}(dep))\cap\mathbf{Feat}(\mathcal{P}(rep))|}{|\mathbf{Feat}(\mathcal{P}(dep))\cup\mathbf{Feat}(\mathcal{P}(rep))|}$. The overall contextual similarity between $\mathcal{O}$ and $\mathcal{U}$ was determined by averaging the cxt-sim scores of paired prompt groups across all API mappings. The final similarity is 0.266. This low similarity highlights a significant difference in the contextual characteristics between the prompts from $\mathcal{O}$ and $\mathcal{U}$, influencing the model’s reference decoding direction (cf. Section IV-A2).
@@
Summary 6: The contextual characteristics of the input prompts, such as the defined variables and function calls, contribute a significant influence to the deprecated API usage.
IV-C RQ3: Library Perspective for Status Quo and Root Causes
@@
We conducted a detailed analysis to examine the LLMs’ completions across different libraries.
@@ IV-C1 Status Quo Analysis
We observe how the AUP and DUR metrics vary with different libraries.
@@
Summary 7: Most LLMs exhibited relatively low API usage plausibility across the 8 libraries, with AUP below 30%. CodeLlama-7b, CodeGen-6b, and CodeGen-2b achieved better AUP for Pandas (about 45%-65%) and TensorFlow (around 40%).
Deprecated Usage Rate. The results are presented in the scatter plots shown in Figure 5, where each data point represents the DUR of a particular LLM for a specific library. The results reveal significant differences in the usage of deprecated APIs across the libraries, with DUR ranging from approximately 0% to 100%. Specifically, LLMs generally showed low DUR (around or below 20%) for Numpy, scikit-learn, and TensorFlow. In contrast, LLMs exhibited consistently high DUR for SciPy (approximately 30%-50%) and PyTorch (approximately 50%-70%), and unstable DUR for Pandas (approximately 30%-80%), seaborn (approximately 20%-60%), and Transformers (approximately 40%-100%).
@@
Summary 8: LLMs showed significant differences in the usage of deprecated APIs across various libraries, with DUR ranging from approximately 0% to 100%. LLMs exhibited consistently high DUR for SciPy and PyTorch, and unstable DUR for Pandas and Transformers.
IV-C2 Potential Cause Discussion
@@
The characteristics of API deprecations during library evolution also significantly impacted the use of deprecated APIs in LLM-based code completions. Some APIs were deprecated due to simple package refactoring, leading to similar usage patterns for the deprecated APIs and their replacements. For example, in the version 1.9.0 release of PyTorch, many APIs for tensor linear algebra were moved from the torch package to the torch.linalg package without changes to the API parameters or usage patterns (e.g., torch.lstsq $\to$ torch.linalg.lstsq). In such cases, LLMs found it more difficult to distinguish between deprecated APIs and their replacements, resulting in more frequent use of deprecated APIs in the predicted completions. To further validate this attribution, we gathered prompts associated with each library from $\mathcal{O}$ and $\mathcal{U}$, denoted as $\mathcal{O}^{l}$ and $\mathcal{U}^{l}$. We then performed feature extraction and calculated the contextual similarity between $\mathcal{O}^{l}$ and $\mathcal{U}^{l}$ following the same approach as in Section IV-B2. A higher contextual similarity indicates that $\mathcal{O}^{l}$ and $\mathcal{U}^{l}$, before and after API deprecation, share more common contextual characteristics and usage patterns. Higher $\mathcal{O}^{l}$-$\mathcal{U}^{l}$ contextual similarity suggests that it is more challenging for LLMs to discern whether to use deprecated or replacing APIs based on the given prompts. Table IV shows the relationship between $\mathcal{O}^{l}$-$\mathcal{U}^{l}$ contextual similarity and DUR ranges across the eight libraries, indicating that the DUR range generally increases with higher $\mathcal{O}^{l}$-$\mathcal{U}^{l}$ contextual similarity.
TABLE IV: Relationship between $\mathcal{O}^{l}$-$\mathcal{U}^{l}$ Contextual Similarity and DUR Range across Eight Libraries
  Library Similarity DUR Range
  Numpy 0.056 $<$ 20%
Pandas 0.193 30%-80%
scikit-learn 0.060 $<$ 20%
SciPy 0.094 30%-50%
seaborn 0.054 20%-60%
TensorFlow 0.078 $<$ 25%
PyTorch 0.126 50%-70%
Transformers 0.672 40%-100%
Summary 9: The characteristics of API deprecations during library evolution significantly impacted the use of deprecated APIs in LLM-based code completions. Minor changes between deprecated APIs and their replacements, such as simple package refactoring, often led to more pronounced issues with deprecated API usages.
V Lightweight Mitigations
@@ V-A Motivation
During model training, a direct mitigation is to clean up the code containing deprecated APIs from the training corpora. However, this is impractical due to the constant library evolution. New API deprecations would cause repeatedly training corpora rebuilding and model retraining. From the library perspective, we cannot control the evolution of the library and API deprecation. Therefore, mitigation should focus on decoding strategies in model inference and prompt engineering.
@@ V-B Two Lightweight Approaches
where $\mathbb{T}^{I}$ and $\mathbb{T}^{O}$ are the input token sequence and output token sequence, respectively. In the context of LLM-based code completion, $\mathbb{T}^{I}$ corresponds to the input prompt $pmpt$, and $\mathbb{T}^{O}$ corresponds to the predicted completion $comp$ (line 1). When $comp$ contains a deprecated API $dep$ (line 3), we need to perform fixing approaches, i.e., the Fix procedure, to replace $dep$ with the corresponding replacement $rep$ (line 4). Note that during the Contains procedure in line 3, alias resolution is conducted similarly to the process in Section III-B1.
@@ 4             $comp\leftarrow\textbf{{Fix}}(pmpt,comp,dep,rep)$
The Fix procedure includes reconstructing input, through either (i) replacing the deprecated API tokens or (ii) inserting additional replacing prompts, and then regenerating output:
  • Approach 1 - ReplaceAPI: Replacing Deprecated API Tokens then Regenerating. As shown in Algorithm 2, ReplaceDep removes the tokens corresponding to $dep$ and any subsequent tokens from $comp$, and then appends the tokens of $rep$ (line 2). This results in a prefix $prefix$ that includes $rep$. The $prefix$ is concatenated with the $pmpt$, and the LLM generates the suffix (line 3), i.e., arguments for $rep$ and the remaining tokens to complete the code line. This concatenation forms the fixed completion $comp^{*}$.
  • Approach 2 - InsertPrompt: Inserting Additional Replacing Prompt then Regenerating. As shown in Algorithm 3, CreateDepPmpt constructs an additional replacing prompt $pmpt^{\prime}$ (line 2), formatted as inline comments to guide the LLM to use $rep$ instead of $dep$. By expressing the replacing instruction into inline comments, the re-written prompt (i.e., $pmpt\oplus pmpt^{\prime}$) can be naturally processed by the LLM to continue writing the code. The replacing prompt is structured as follows, where “{comp}”, “{dep}”, and “{rep}” are placeholders for the original completion, deprecated API, and replacing API, respectively, and “...” represents the indentation to ensure syntax correctness.
@@ 4       $comp^{*}\leftarrow prefix\oplus suffix$
Algorithm 2 Approach 1 - ReplaceAPI
@@ 3       $comp^{*}\leftarrow\textbf{{LLM}}(pmpt\oplus pmpt^{\prime})$
Algorithm 3 Approach 2 - InsertPrompt
@@ V-C1 Evaluation Procedure
For each LLM, we selected up-to-dated samples from $\mathcal{U}$ where the LLM predicted DepC completions using deprecated APIs, i.e.,
$\mathcal{T}=\{(pmpt,dep\to rep)\in\mathcal{U}:\text{anno}(\textbf{{LLM}}(pmpt))=\emph{DepC}\},$
to constitute the evaluation data. These samples were chosen because they have corresponding ground-truth completions (e.g., the line following the prompt in Figure 2), which are essential for assessing the effectiveness of the proposed fixing approaches. For each sample, we employed the deprecation-aware code completion illustrated in Algorithm 1 with the two fixing approaches ReplaceAPI and InsertPrompt to prompt the LLM to generate a completion. We used the following three metrics to assess their effectiveness:
  • Fixed Rate (FR): This metric indicates the proportion of RepC completions predicted by the fixing approaches.
@@ V-C2 Results and Analysis
The evaluation results of the proposed fixing approaches are presented in Table V.
TABLE V: Evaluation Results of Proposed Approaches
@@
ReplaceAPI. Using the ReplaceAPI fixing approach (Algorithm 2), all the LLMs achieve high fixed rates (FR), with values exceeding 85%. Failures in fixing are mainly due to syntax errors or incorrect function calls caused by erroneous tokens following the replaced APIs. For example, consider a DepC completion: “meta_graph_def = tf.saved_model.loader.load(…)” predicted by the original completion procedure (Algorithm 1, line 1). ReplaceAPI replaces the deprecated API tf.saved_model.loader.load with its replacement tf.saved_model.load, producing a $prefix$ of “meta_graph_def=tf.saved_model.load” (line 2 of Algorithm 2). However, CodeGen-2b then predicts a $suffix$ of “_meta_graph_def(…)” (line 3 of Algorithm 2), resulting in an erroneous function call: tf.saved_model.load_meta_graph_def(). This issue arises because the replacement operation in ReplaceAPI can disrupt the naturalness of the code context [75] and the LLMs’ decoding process. An additional interesting finding is that for the three versions of CodeGen, the FR decreases as the model size increases, suggesting that larger models might be more sensitive to interventions during decoding.
The completions generated by LLMs using the ReplaceAPI approach exhibit high ES of over 80%, and EM rates between 30% and 50%, with these rates increasing alongside model size. The inaccuracies in the completions often involve incorrect return values and arguments for the replacing APIs. Incorrect return values arise because the replacing API might include different elements compared to the deprecated API, and the ReplaceAPI approach cannot resolve such inconsistencies in the $prefix$ (line 2 of Algorithm 2). The incorrect arguments are primarily due to the LLMs’ limitations in correctly utilizing replacing APIs, especially those with complex argument lists.
InsertPrompt. When applying the InsertPrompt fixing approach (Algorithm 3), the LLMs exhibited significantly varied fixed rates, ranging from 25.7% to 97.2%. This variation suggests that larger models generally possess a stronger capacity to interpret and utilize inserted prompts formatted as inline comments. An exception to this trend is DeepSeek-1.3b, which achieved a notably high FR of 93.5%. This success can be attributed to using the deepseek-coder-1.3b-instruct version, which has robust zero-shot instruction-following capabilities. Moreover, the FR differences among various LLMs highlight their sensitivity to prompt construction [76, 77], which indicates that different LLMs may require specialized additional prompts in InsertPrompt for optimal performance.
Summary 10: ReplaceAPI effectively addresses deprecated API usage for all open-source LLMs, achieving fix rates exceeding 85%. The fixed completions demonstrate acceptable accuracy. While InsertPrompt does not achieve sufficient effectiveness and accuracy in fixing completions containing deprecated API usage, it provides future potential of exploration.
@@ VI-A Implications
Validating LLM-Generated Code Completions and Issuing Alerts for Deprecated API Usages. Our evaluation study reveals that LLMs frequently use deprecated APIs during code completion. Such deprecated API usages can be easily overlooked [25], potentially introducing bugs or security vulnerabilities into software projects. Therefore, implementing a validation mechanism for deprecated API usage in LLM-generated code completions is crucial to ensure the reliability of the code. Such a validation mechanism can be further integrated into the post-processing of LLM-based code completion, such as issuing alerts to developers.
@@
Designing Agent & Multi-Agent Systems for Incorporating Library Evolution into LLM-Driven Software Development. In modern software development driven by LLMs, the issues brought by library evolution are encountered not only in code completion. With advancements in AI agents [87, 88, 89], we can potentially develop autonomous agents or multi-agent systems capable of automatically discovering deprecated API usages, identifying correct replacements, upgrading dependent libraries, and fixing the code [90]. To ensure comprehensive recognition of deprecated API usage and up-to-date fixes, we should design an effective multi-agent collaboration pipeline. One agent should scan the generated code to identify all pieces related to API usage. Another agent should continuously fetch information online, checking the latest official API documentation to aid in discovery and correction. Finally, a dedicated agent should be responsible for implementing the necessary library upgrades and code fixes.
@@ VI-B Threats to Validity
Internal Threats. The primary threat to our study is the soundness of the static analysis used for function location and result annotation. Given that Python is a dynamic programming language, the lightweight object type resolution and alis resolution we employed may have missed some function calls of deprecated and replacing APIs during the matching process. Nevertheless, they did not contribute bias to deprecated APIs or replacing APIs in our extracted function calls. In the future, we aim to address this issue by implementing advanced type inference techniques. Additionally, our study currently focuses on function-level API deprecation, overlooking parameter-level deprecations. Future research would investigate a broader type of deprecated APIs to provide more comprehensive analysis.
External Threats. A primary threat lies in the choice of Python libraries and LLMs. To mitigate this threat, we reused libraries examined in previous studies and introduced three popular deep learning libraries to ensure diversity and timeliness. For the LLMs, we selected models covering various architectures, model sizes, training corpora, and training strategies to ensure the generalizability of our findings. Another external threat is that the study was conducted solely on the Python language, which may limit the applicability of our findings to other languages such as Java and C#. In the future, we plan to explore the impact of library evolution on LLM-based code completion across a broader range of programming languages.